Compute the value of the discriminant and give the number of real solutions to the quadratic equation.
-5x^2 -2x+3=0
Please help!!!!!!! and Thank You!!!!!!

Answer by mukhopadhyay(490)

Discriminant for quadratic equation ax^2+bx+c =0 is (b^2-4ac)
Here a=-5, b=-2 and c=3
(b^2-4ac) = 4-4(-5)(3) = 4+60 = 64
The real solution to this equation are:
-b/2a + sqrt(discriminant)/2a and -b/2a - sqrt(discriminant)/2a
Same as, x= -(-2)/(-10) + 8/(-10) = -2/10 - 8/10 = -1
and x= -2/10 + 8/10 = 6/10 = 3/5
Answer: x=-1 or x=3/5